In this class, one of your 4 test scores is dropped. On the first three tests, you have the following scores: 70, 31, and 95. What score do you need on the next test to raise your average (incorporating the dropped test) to 75?
Answer #1

= X / n

75 = (70 + 31 + 95 + x ) / 4

75 = ( x + 196) / 4

Solve for x

x = 104